A train leaves town at 7:00am traveling south at 30mph. At 10:00 that morning, a second train leaves from the same location (on a parallel track) traveling 40 mph. At what time will it overtake the slower? You don't have to solve the problem for me. I just need help setting this up.
Answer by stanbon(75887)

A train leaves town at 7:00am traveling south at 30mph. At 10:00 that morning, a second train leaves from the same location (on a parallel track) traveling 40 mph. At what time will it overtake the slower? You don't have to solve the problem for me.
--------------------------
Early train DATA:
rate = 30 mph ; time = x hrs ; distance = 30x miles
----------
Later train DATA:
rate = 40 mph ; time = x-3 hrs ; distance = 40(x-3) miles
-----
Equation:
distance = distance
30x = 40(x-3)
==============================
Cheers,
Stan H.
==============================